2015, equipment purchased for 42750. 4 years residual value 3350. 2017 useful life determined is 9 years (another 7 years) with residual value 7500. using straight line method what is deprec. in 2017
Answer #1

Depreciation of 2017= $ 2,221.43 or 2,221

Working

Straight line Method
A Cost $ 42,750
B Residual Value $ 3,350
C=A - B Depreciable base $ 39,400
D Life [in years left ] 4
E=C/D Annual SLM depreciation $ 9,850.00

.

Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
2015 $ 42,750.00 $ 9,850.00 $ 9,850.00 $ 32,900.00
2016 $ 32,900.00 $ 9,850.00 $ 19,700.00 $ 23,050.00

.

Straight line Method
A Ending value at year 2016 $ 23,050
B Residual Value $ 7,500
C=A - B Depreciable base $ 15,550
D Life [in years left ] [9-2] 7
E=C/D Annual SLM depreciation $ 2,221.43